Roadside market has 8.45 percent coupon bonds outstanding that mature in 10.5 years. The bonds pay interest semiannually. What is the market price per bond if the face value is $1000 and the yield to maturity is 7.2 percent ? show with steps and formulas. no excel or word please
Answer #1

(In $)

Market value of bond

= Coupon (Present value of annuity @ Ytm, n) + Redemption value (Present value @Ytm, n)

Since the bond is compounding semi annually

= 1000×8.45% (1/2) × (PVAF @ 3.6%, 21 periods) + 1000(PVF @ 3.6%, 21st period)

*Assume Redemption value is equals to Face value

= 42.25(14.560) + 1000(0.476)

= 615.16 + 476

= $ 1091.16